Required information [The following information applies to the questions displayed below.j In 2018, the Westgate Construction Company entered into a contract to construct a road for Santa Clara County for $10,000,000. The road was completed in 2020. Information related to the contract is as follows 2019 $2,490,000 $3,984,000 $2,008,600 1,826,000 2,030,000 4,444,000 3,526,000 3,900,000 2018 5,810,000 1,815,000 2020 Cost incurred during the year Estimated costs to complete as of year-end Billings during the year Cash collections during the year 4,285,000 Westgate recognizes revenue over time according to percentage of completion Requirec 1. Calculate the amount of revenue and gross profit (loss) to be recognized in each of the three years. (Do not round intermediate calculations. Loss amounts should be indicated with a minus sign.) 2018 2019 2020 Revenue Gross profit (loss)
Answer #1

Answer

  • All working forms part of the answer
  • Please follow the following schedule of working

Working

2018

2019

2020

A

Contract Price

$     10,000,000.00

$    10,000,000.00

$        10,000,000.00

B

Cost Incurred to Date

$       2,490,000.00

$       6,474,000.00

$          8,482,600.00

C

Estimated cost yet to be incurred to complete the contract

$       5,810,000.00

$       1,826,000.00

$                               -  

D = B+C

Total Cost

$       8,300,000.00

$       8,300,000.00

$          8,482,600.00

E = (B/D) x 100

% of Completion

30.00%

78.00%

100.00%

F = A x E

Revenue to date

$       3,000,000.00

$       7,800,000.00

$        10,000,000.00

G

Revenue of Previous year

$                            -  

$       3,000,000.00

$          7,800,000.00

H = F - G

Net Revenue this year

$       3,000,000.00

$       4,800,000.00

$          2,200,000.00

I (=B)

Cost to date

$       2,490,000.00

$       6,474,000.00

$          8,482,600.00

J

Cost to date of previous year

$                            -  

$       2,490,000.00

$          6,474,000.00

K = I - J

Net Cost for the year

$       2,490,000.00

$       3,984,000.00

$          2,008,600.00

L = H - K

Gross Profits

$           510,000.00

$          816,000.00

$              191,400.00

  • Answers

2018

2019

2020

Revenues

$       3,000,000.00

$       4,800,000.00

$          2,200,000.00

Gross Profits

$           510,000.00

$          816,000.00

$              191,400.00